A man leaves home driving at the rate of 50mph. When his daughter discovers that he has forgotten his wallet, she drives after him at the rate of 65mph. How long will it take her to catch her dad if he has a 15-minute head start?
Answer by mananth(16949)

Man's speed 50 mph
Daughter's speed 65 mph
Difference in time= 00:15 => 0.25 hours
0.25*50= 12.5
Man will have covered 12.5 miles before daughter starts
catch up distance= 12.5 miles
catch up speed = 65 -50 mph
catch up speed = 15 mph
Catchup time = catchup distance/catch up speed
catch up time= 12.5 / 15
catch up time= 0.83 hours => 49.8 minutes
50 minutes